# Going Forward-Forward in Distributed Deep Learning

### v1

@@ 5.2 Classifier Model Comparison
Adaptive-NEG Goodness Single Layer 5,254.87 98.43
All Layers 2,980.76 98.51
@@
Adaptive-NEG Softmax Single Layer 2,471.27 98.31
All Layers 1,886.42 98.30
5.3 Softmax Comparison of Random-NEG and Fixed-NEG
In this section, we compare the Softmax method with Fixed-NEG and Random-NEG, with results given in the Table 3.
We first observe that the accuracy of the Softmax approach is slightly higher than the goodness prediction.
@@ 5.3 Softmax Comparison of Random-NEG and Fixed-NEG
As expected, training a model with additional Softmax layer takes longer in the Sequential model. However, Single Layer implementation is not effected from this additional complexity. This is due to the fact that Softmax Layer is easier/faster to train than the FF Layers and thanks to the pipeline architecture this only adds a small delay.
@@
Surprisingly, Softmax models give faster results compared to same Goodness models for All Layers implementation. This again stems from the fact that training Softmax is faster than training FF Layers and distributing this job between nodes speeds ups the overall training.
Table 3: Softmax Comparison of Random-NEG and Fixed-NEG
Model Implementation Time (s) Goodness Softmax
Accuracy (%) Accuracy (%)
Sequential 8,104.96 98.33 98.48
Random-NEG Softmax Single Layer 1,891.86 98.20 98.31
All Layers 1,786.30 98.16 98.33
Sequential 7,178.71 98.33 -
Random-NEG Goodness Single Layer 1,974.15 98.26 -
All Layers 2,008.25 98.17 -
Sequential 7,983.21 97.95 98.03
Fixed-NEG Softmax Single Layer 1,939.62 98.01 97.92
All Layers 1,794.52 97.86 97.85
Sequential 7,143.28 97.95 -
Fixed-NEG Goodness Single Layer 1,920.81 97.94 -
All Layers 1,978.21 97.89 -
@@ 6 Conclusion and Future Work
Our work in this paper presents Pipeline Forward-Forward Algorithm (PFF), a novel way of training distributed neural networks using Forward-Forward Algorithm. Compared to the classic implementations with backpropagation and pipeline parallelism [2] [3], PFF is inherently different as it does not enforce the dependencies of backpropagation to the system, thus achieving higher utilization of computational units with less bubbles and idle time. Experiments that are done with PFF shows that PFF Algorithm achieves the same accuracy as standard FF implementation [5] with a 4x speed up. Comparison of PFF with an existing distributed implementation of Forward-Forward (DFF [10]) shows even greater improvements as PFF achieves %5 more accuracy in 10 times less epochs. In addition, the data that is exchanged between layers in PFF is a lot less than DFF since PFF sends the layer information (weights and biases) whereas DFF sends the whole output of the data. This results in less communication overhead compared to DFF.

### v2

@@ 5.3 Classifier Mode Comparison for AdaptiveNEG
AdaptiveNEG-Goodness Single-Layer 5,254.87 98.43
All-Layers 2,980.76 98.51
@@
AdaptiveNEG-Softmax Single-Layer 2,471.27 98.31
All-Layers 1,886.42 98.30
5.4 Classifier Mode Comparison for RandomNEG
Building upon previous findings in the section 5.2, we observed that RandomNEG closely rivals AdaptiveNEG in accuracy, while significantly reducing training time. This observation leads us to further investigate the impact of employing softmax within the RandomNEG framework. This section thus compares the effects of softmax and Goodness prediction on both FixedNEG and RandomNEG strategies. The results of this comparison are detailed in Table 3, offering insights into how these elements influence the effectiveness and efficiency of the methods under review.
In terms of accuracy, the Softmax approach outperforms the Goodness approach for prediction. For instance, while Sequential RandomNEG Goodness model performed 98.33 the softmax version performed 98.48. This accuracy level of the softmax version is very close to the performance of our baseline model (98.51 Accuracy using AdaptiveNEG Goodness). While the accuracy levels are very close we observe that there is a significant speed up using RandomNEG (8,104 versus 11,190).
@@ 5.4 Classifier Mode Comparison for RandomNEG
In terms of training speed on the other hand, training a model with the additional Softmax layer takes longer in the Sequential model as expected. However, Single-Layer implementation is not effected from this additional complexity. This is due to the fact that Softmax Layer is easier/faster to train than the FF Layers and thanks to the pipeline architecture this only adds a small delay.
@@
Surprisingly, Softmax models give faster results compared to same Goodness models for All-Layers implementation. This again stems from the fact that training Softmax is faster than training FF Layers and distributing this job between 5 nodes speeds ups the overall training.
Table 3: Classification mode comparison for RandomNEG. The first three row results indicated in italic are from Table 1. Bold text shows the proposed RandomNEG model balancing accuracy and speed.
Model Implementation Training Time (s) Test Accuracy (%)
Sequential 7,178.71 98.33
RandomNEG-Goodness Single-Layer 1,974.15 98.26
All-Layers 2,008.25 98.17
Sequential 8,104.96 98.48
RandomNEG-Softmax Single-Layer 1,891.86 98.31
All-Layers 1,786.30 98.33
5.5 Evaluation of the Performance-Optimized Model
This section explores the evaluation of our Performance-Optimized Model that is created by incorporating a new goodness function into the Forward-Forward algorithm. Our analysis shows that while the Performance-Optimized Model performs slightly worse than the best-performing AdaptiveNEG Goodness model, there is a notable reduction in training times, indicating a significant step towards efficient computing without significantly sacrificing performance.
In the MNIST dataset, the results presented in Table 4 highlight the effectiveness of the proposed new goodness function when integrated with the Forward-Forward algorithm. The highest test accuracy is achieved by the AdaptiveNEG Goodness model, reaching 98.52%, which is slightly superior to the 98.48% accuracy of the RandomNEG Softmax model. It’s noteworthy that the Performance-Optimized Network, whether using only the last layer or all layers, shows a small decrease in accuracy. Specifically, using all layers improves accuracy slightly to 98.38% compared to 98.30% when only the last layer is used. These results underscore the robustness of the proposed goodness function, particularly when all layers are optimized, despite a considerable reduction in training time (4219.97 seconds) compared to the AdaptiveNEG Goodness model (11190.72 seconds).
Table 4: MNIST evaluation of Performance-Optimized Model with the baseline models from Table 1.
Model Training Time (s) Test Accuracy (%)
AdaptiveNEG-Goodness 11,190.72 98.52
RandomNEG-Softmax 8,104.96 98.48
Performance-Optimized Network (only last layer) 4,219.97 98.30
Performance-Optimized Network (using all layers) 4,219.97 98.38
5.6 Experiments with CIFAR-10
To extend our analysis to a more complex scenario, we replicated the experiments described in sections 5.2 through 5.5 using the CIFAR-10 dataset. CIFAR-10, known for its higher variability and complexity compared to MNIST, serves as a challenging benchmark for testing the robustness and effectiveness of our models.
Table 5 presents the detailed results of these experiments. The Performance-Optimized Network, whether using only the last layer or utilizing all layers, demonstrated a superior ability to handle the complexities of CIFAR-10, with the all-layers implementation achieving the highest accuracy of 53.50%. This was closely followed by the RandomNEG Softmax model, which achieved 52.18% accuracy. Furthermore, the Performance-Optimized Network models shows relatively efficient training durations compared to other PFF Models. Surprisingly, the AdaptiveNEG Goodness model, which performed exceptionally well on MNIST, significantly underperformed on CIFAR-10, with an accuracy of only 11.10%. This difference suggests a potential mismatch between the model’s capabilities and the dataset’s requirements.
It should be noted that while state-of-the-art models have achieved beyond 99% accuracy on CIFAR-10, the performance of our models are similar to Hinton’s original FF Algorithm, which reported an accuracy of 56% using CNNs [5].
Table 5: CIFAR-10 Results
Model Training Time (s) Test Accuracy (%)
Performance-Optimized Network (using all layers) 4,920.97 53.50
Performance-Optimized Network (only last layer) 4,920.97 53.11
FixedNEG-Softmax 8,021.15 50.89
RandomNEG-Softmax 7,636.99 52.18
AdaptiveNEG-Goodness 10,148.23 11.10
@@ 6 Conclusion and Future Work
Our work in this paper presents Pipeline Forward-Forward Algorithm (PFF), a novel way of training distributed neural networks using Forward-Forward Algorithm. Compared to the classic implementations with backpropagation and pipeline parallelism [2] [3], PFF is inherently different as it does not enforce the dependencies of backpropagation to the system, thus achieving higher utilization of computational units with less bubbles and idle time. Experiments that are done with PFF shows that PFF Algorithm achieves the same accuracy as standard FF implementation [5] with a 4x speed up. Comparison of PFF with an existing distributed implementation of Forward-Forward (DFF [11]) shows even greater improvements as PFF achieves %5 more accuracy in 10 times less epochs. This improvement in accuracy is mainly because PFF splits the data into batches and feeds it to the network batch by batch unlike DFF which feeds the data as whole. In addition, the data that is exchanged between layers in PFF is a lot less than DFF since PFF sends the layer information (weights and biases) whereas DFF sends the whole output of the data. This results in less communication overhead compared to DFF.